OpenStudy (anonymous):

Which of the following represents the area of a rectangle whose length is 3x − 7 and whose width is x + 2?

OpenStudy (anonymous):

3x2 − 14 3x2 + x − 14 3x2 − x − 14 3x2 + 14

OpenStudy (anonymous):

(3x−7x)+2 =3x+(−7x)+2 =(−4x)+2 =−4x+2

OpenStudy (anonymous):

C

OpenStudy (anonymous):

I mean the last one

OpenStudy (anonymous):

?